OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5

OpenStudy (anonymous):

can some one help?

OpenStudy (mrhoola):

Are you familiar with the three methods to solve linear equations ? Substitution , Addition , I forgot what the other one was

OpenStudy (anonymous):

yes a little bit

OpenStudy (kropot72):

You can use to first equation to find y in terms of x: x - y = -3 Subtracting x from both sides we get -y = -x - 3 Multiplying both sides by -1 gives us y = x + 3 Now plug the value for y into the second of the original equations, giving: x + 3(x + 3) = 5 ...................(1) Now you just need to solve equation (1) to find the value of x.

OpenStudy (anonymous):

how do i solve for 1?

OpenStudy (anonymous):

well how do i solve x+3(x+3)=5

OpenStudy (kropot72):

x + 3(x + 3) = 5 ...................(1) Expanding the brackets gives x + 3x + 9 = 5 Adding like terms we get 4x + 9 = 5 ............(2) Now you should be able to solve equation (2).

OpenStudy (anonymous):

4x+9=5 -9 -9 4x=-4

OpenStudy (anonymous):

is that right so far?

OpenStudy (kropot72):

Yes, 4x = -4 is correct. Now divide both sides by 4.

OpenStudy (anonymous):

x=-1?

OpenStudy (kropot72):

You are correct.

OpenStudy (anonymous):

alright thanks man, greatly appreciate your help

OpenStudy (kropot72):

You're welcome :)